Keith borrowed money from a credit union for 5

 years and was charged simple interest at an annual rate of 7

%

. The total interest that he paid was $

3150

. How much money did he borrow?


1
Expert's answer
2022-02-10T13:09:20-0500

"Interest, I=\frac{PRT}{100}"

"3150=\frac{P\times 7 \times 5}{100}"

"P=\frac{3150 \times 100}{35}=\$9000"